OpenStudy (anonymous):

A rectangle has an unknown width and length. The length is 4 less than twice the width. The area of the rectangle is 16. What is the width and length of the rectangle?

OpenStudy (anonymous):

Need help!!

OpenStudy (jdoe0001):

w = width l = length "length is 4 less than twice the width" "twice the width" => 2w "4 less than twice the width" => 4- 2w  \(\bf w = width\qquad \qquad l = length=4-2w\\ \quad \\ Area = 16 = l\times w\implies (4-2w)\times w\\ \quad \\ \quad \\ 16 = (4-2w)\times w\) solve for "w" to know what the width :) whatever the width will be, the length will be, 4 -2w

OpenStudy (jdoe0001):

hmmm.... actually hold on, 4 less than twice will be 2w-4 so lemme rewrite that quick :)

OpenStudy (jdoe0001):

 \(\bf w = width\qquad \qquad l = length=2w-4\\ \quad \\ Area = 16 = l\times w\implies (2w-4)\times w\\ \quad \\ \quad \\ 16 = (2w-4)\times w\) solve for "w" to know what the width :) whatever the width will be, the length will be, 2w-4